OpenStudy (anonymous):

Which of the following represents 3x - 5y + 10 = 0 written in slope-intercept form? y = -3/5x + 2 y =3/5 x - 2 y =3/5 x + 2

OpenStudy (anonymous):

lope-intercept form y = mx + c

OpenStudy (anonymous):

Change 3x - 5y + 10 = 0 this into the form y = mx + c

OpenStudy (anonymous):

3x - 5y + 10 = 0 -5y = -10 - 3x 5 y = 3x + 10 y = 3x/5 + 2